Duncan takes a break from studying and goes to the gym to swim laps. If swimming burns 6.95×105 cal per hour, how many kilojoules does swimming burn in that same amount of time? In other words, convert 6.95×105 cal to kilojoules. Express your answer numerically in kilojoules.
Answer #1

The value of 1 cal = 4.184 J

6.95 * 105 cal = 6.95 * 105 * 4.184 J = 2.91 * 103 KJ